# Monte Carlo Tree Diffusion for System 2 Planning

### v3

@@ 5.2 Maze Navigation with Point‐Mass and Ant Robots
Results.
Table 1 presents success rates across medium, large, and giant mazes for both point‐mass and ant robots. We can see that MCTD consistently surpasses other methods by a large margin. Notably, Diffuser‐Random Search, despite using roughly the same number of denoising steps as MCTD, demonstrates no improvement over one‐shot Diffuser, underscoring the importance of tree-based systematic branching over random sampling. Diffuser‐Replanning similarly fails to outperform the base Diffuser, whereas Diffusion Forcing exhibits higher success rates in larger mazes, indicating benefits from a semi‐autoregressive schedule for long horizons. However, only MCTD achieves near‐perfect performance on the most challenging (large, giant) mazes, thanks to its ability to refine partial trajectories and avoid dead‐ends or inconsistent plan segments. This process is visualized in Figure 3. The performance on the giant map is impressive, because in the map, there are lots of detour pathes to reach the goal as shown in Figure 2. Comparing with Diffuser‐Random Search further highlights MCTD’s efficiency, since it offers substantially better outcomes under the same computational budget.
In the antmaze domain, MCTD continues to perform nearly perfectly, while baseline performance generally degrades more significantly than in pointmaze, except for the medium-sized settings under Diffusion Forcing. A principal reason for the baseline’s degradation is that the value learning policy (dql) struggles in high‐dimensional control tasks. Even the performance degradations exist due to inaccurate execution, MCTD consistently outperforms baselines and the gap is larger as the maze size is larger.
@@ 5.3 Robot Arm Cube Manipulation
For multi‐cube manipulation tasks from OGBench, a robot arm must move one to four cubes to specific table locations as shown in Figure 4(a). Increasing the number of cubes grows both the planning horizon and the combinatorial complexity. While the planner (MCTD or a baseline) issues high‐level positions for the cubes, a value-learning policy (dql) executes local actions. We augment MCTD with object‐wise guidance by selecting which cube is manipulated at each step, avoiding simultaneous movements of multiple cubes. We note that the object-wise guidance is particularly suited to MCTD because the tree-expansion mechanism naturally supports discrete ”meta-actions” at each node, such as selecting which object to move next. Single-pass diffusion methods, by contrast, attempt to produce a single, holistic trajectory in one go, leaving no opportunity to decide object-by-object. This guidance and additional task‐specific reward shaping are described in Appendix A.6.
((a))
((b))
((c))
Figure 4: (a) Illustration of the robot arm cube manipulation task. (b) A holistic plan for moving two cubes, where entanglement between their partial trajectories leads to an inaccurate movement (blue cube). (c) An object‐wise plan produced by MCTD‐Replanning, which avoids entanglement and achieves more reliable control over each cube’s movement.
Results.
Table 2 shows success rates for one to four cubes. All methods perform comparably on single‐cube tasks, but performance drops significantly for multi‐cube scenarios, especially for standard diffusion baselines that struggle to decide cube ordering and handle subtasks such as stacking. However, MCTD-Replanning shows significant performance gaps from other models for multi-cube scenarios. MCTD exhibits moderate gains (e.g., 22% on two cubes), though it suffers from holistic plan entanglements when multiple objects are involved as shown in Figure 4(b). To address this, MCTD‐Replanning periodically replans, effectively separating each cube’s movements as shown in Figure 4(c). This boosts success on the double‐cube problem from 22% to 50%, and also benefits tasks with more cubes. Another challenge arises from the value-learning policy, which generalizes poorly as the number of cubes increases. Even if MCTD generates a suitable global plan, flawed local control undermines the final outcome. Interestingly, training the policy solely on single‐cube data still allows MCTD‐Replanning to solve a portion (24%) of quadruple‐cube scenarios, illustrating the potential of object‐wise guidance and iterative replanning.
@@ 5.4 Visual Pointmaze
To evaluate our method in image-based planning, particularly under partial observability, we introduce a visual pointmaze environment where the agent receives top-down pixel images of the initial and goal states, as shown in Figure 5. Unlike the visual antmaze in OGBench, which involves complex locomotion, this environment allows us to focus on how well the planner operates with raw visual inputs without relying on ground-truth states.
@@
We pre‐encode observations with a Variational Autoencoder (VAE) (vae) to generate latent states suitable for diffusion‐based planning. Because Diffuser’s default heuristic controller is inapplicable to purely visual inputs, we design an inverse dynamics model that maps consecutive latent observations to actions, $\hat{a}_{t}=\mathrm{InvDyna}\bigl{(}(x_{t-1},\,x_{t}),\,x_{t+1}\bigr{)}$. For guidance, we employ a position estimator trained on ground‐truth coordinates; however, this estimator does not update the underlying VAE, ensuring that the planner remains in a partially observable regime. The details for this task setting are discussed in Appendix A.7.
Results.
@@
Table 3 compares success rates on medium and large versions of the visual pointmaze. MCTD and MCTD-Replanning show better performance than baselines on image-based, partial observable tasks. MCTD outperforms Diffuser baselines and Diffusion Forcing in the medium maze, presumably due to its tree‐structured exploration, which better balances exploration and exploitation under perceptional partial observability. In the large maze, single‐pass planners fail almost entirely, highlighting the difficulty of long‐horizon visual planning. MCTD‐Replanning improves performance by reevaluating partial plans mid‐trajectory, though success remains modest overall. This result underscores the inherent challenge of scaling diffusion‐based methods to extended horizons with partial observational visual data and suggests that future work may need more sophisticated visual encoders or hierarchical guidance strategies.
@@
Figure 6: Success rates (%) and wall‐clock runtime (Sec.) as the maximum number of denoising steps (test‐time budget) increases for large and giant pointmaze tasks
5.5 Test-Time Compute Scalability & Time Complexity
@@
We examine how each planner leverages additional test‐time computation by varying the maximum number of denoising steps and measuring both success rate and runtime on the large and giant pointmaze tasks (Figure 6). As the TTC budget grows, MCTD consistently achieves high success rates—ultimately nearing perfection in giant mazes. In contrast, simply increasing denoising steps (Diffuser) or drawing more random samples (Diffuser‐Random Search) yields modest returns, suggesting limited test‐time scalability.
@@
MCTD incurs additional overhead from maintaining and expanding a tree—leading to a larger runtime as the budget grows. For instance, Diffuser‐Random Search shows moderate runtime growth because multiple samples can be denoised in parallel batches, which is not directly feasible in MCTD’s tree‐structured approach. Nonetheless, MCTD can terminate early whenever a successful plan is found, thereby limiting unnecessary expansions. As a result, on the large maze, MCTD’s runtime quickly saturates once the task can be solved with a smaller search depth, rather than fully consuming the available budget. A similar pattern emerges in the giant maze: as success rates increase under higher budgets, the runtime growth for MCTD remains sublinear, indicating that many expansions terminate early once suitable trajectories are located.
@@ 6 Limitations & Discussion
While MCTD enhances test-time compute scalability by integrating search-based planning with diffusion models, it remains computationally expensive due to its System 2-style deliberate reasoning. Determining when to engage in structured planning versus relying on fast System 1 (model-free) planning is an open challenge, as expensive tree search may not always be necessary. Adaptive compute allocation based on task complexity or uncertainty could improve efficiency.
Another limitation is inefficiency in large-scale search spaces, where evaluating multiple trajectory hypotheses remains computationally demanding despite using low-dimensional meta-actions. A promising direction is amortizing search learning, where the system meta-learns from test-time search to improve exploration efficiency over time. Instead of treating initial exploration as random, MCTD could incorporate test-time learning mechanisms to refine its search dynamically. Additionally, self-supervised reward shaping could improve trajectory evaluation in sparse-reward settings. Optimizing parallelized denoising, differentiable tree search, and model-based rollouts could further enhance efficiency.
@@ 7 Conclusion
We introduced Monte Carlo Tree Diffusion (MCTD), a framework designed to combine the best of both worlds: the structured search of Monte Carlo Tree Search and the generative flexibility of diffusion planning to enhance the test-time compute scalability of System 2 planning. MCTD leverages meta-actions for adaptive exploration-exploitation, tree-structured denoising for efficient diffusion-based expansion, and fast jumpy denoising for rapid simulation. Experimental results demonstrate that MCTD outperforms existing approaches in various planning tasks, achieving superior scalability and solution quality. Future work will explore adaptive compute allocation, learning-based meta-action selection, and reward shaping to further enhance performance, paving the way for more scalable and flexible System 2 planning.
@@ A.1 Baselines
We compare MCTD against several baselines, each highlighting different ways to leverage test-time computing on diffusion models or broader planning strategies:
  • Diffuser (diffuser): We use the standard single‐shot Diffuser as our primary diffusion baseline. It generates a full trajectory by denoising from maximum to zero noise in one pass, guided by a return‐predicting function $J_{\phi}$. While it can produce coherent plans, it lacks any mechanism to adapt or refine these plans at test time.
  • Diffuser-Replanning: To partially address the lack of iterative search, we evaluated a replanning variant of Diffuser. After an initial denoised plan is sampled, the policy is periodically re‐invoked at predefined intervals (e.g., every 10 time steps), using the agent’s new state as the “start” for a fresh diffusion process. This strategy leverages additional test‐time compute by attempting incremental corrections, but does not share information across different replans.
  • Diffuser-Random Search: We consider a random‐search variant of Diffuser that generates multiple trajectories from the randomly noised plans and picks the best candidate according to a heuristic or learned value function. This approach can be seen as a “Sample-Score-Rank” method proposed in (d-mpc), increasing test‐time compute by drawing more samples, but lacking the systematic exploration or adaptation found in tree‐search paradigms.
  • Diffusion Forcing: Diffusion Forcing (diffusion_forcing) extends Diffuser by introducing a tokenized, causal noise schedule. Subplans of the trajectory can be denoised at different rates, allowing partial re‐sampling of specific time segments without re‐optimizing the entire plan. In our experiments, we evaluate Diffusion Forcing with the Transformer backbone provided in their official repository, which performed better in our preliminary tests. This method helps isolate the benefit of tokenized partial denoising from our tree‐structured approach. We note that Diffusion Forcing baselines are basically designed to do replanning, so we follow the setting in this evaluation.
A.2 Heuristic Controller, Value-Learning Policy and Inverse Dynamics Model
A long‐standing challenge in diffusion‐based planning is the difficulty of preserving global trajectory coherence while managing local, high‐dimensional state‐action control (diffusion_forcing; plandq). For example, PlanDQ (plandq) couples a high‐level diffusion planner with a learned low‐level policy. Similarly, our approach focuses the diffusion planner on lower‐dimensional, representative state information (e.g., object or agent positions), while delegating detailed action inference to the hueristic controller or value-learning policy or inverse dynamics model. The details of this integration is discussed in Algorithm 10. Such a hierarchical design allows MCTD and baseline planners to concentrate on broader strategic decisions while leaving the finer‐grained execution details to the heuristic controller or value-learning policy or inverse dynamics model.

### v5

@@ 5.2 Maze Navigation with Point-Mass and Ant Robots
Results
Table 1 presents success rates across medium, large, and giant mazes for both point-mass and ant robots. MCTD consistently surpasses other methods by a large margin. Notably, Diffuser-Random Search, despite using roughly the same number of denoising steps as MCTD, demonstrates no improvement over one-shot Diffuser, underscoring the importance of tree-based systematic branching over random sampling. Diffuser-Replanning similarly fails to outperform the base Diffuser, whereas Diffusion Forcing exhibits higher success rates in larger mazes, indicating benefits from a semi-autoregressive schedule for long horizons. However, only MCTD achieves near-perfect performance on the most challenging (large, giant) mazes, owing to its ability to refine partial trajectories and avoid dead-ends or inconsistent plan segments. This process is visualized in Figure 3. The performance on the giant map is impressive, because the map contains numerous detour paths to reach the goal as shown in Figure 2. The comparison with Diffuser-Random Search further highlights MCTD’s efficiency, since it offers substantially better outcomes under the same computational budget.
In the antmaze domain, MCTD continues to perform nearly perfectly, while baseline performance generally degrades more significantly than in pointmaze, except for the medium-sized settings under Diffusion Forcing. A principal reason for the degradation in the baselines is that the value learning policy (Wang et al., 2023) struggles in high-dimensional control tasks. Despite performance degradations due to inaccurate execution, MCTD consistently outperforms baselines and the performance gap widens as the maze size increases.
@@ 5.3 Robot Arm Cube Manipulation
For multi-cube manipulation tasks from OGBench, a robot arm must move one to four cubes to specific table locations as shown in Figure 4a. Increasing the number of cubes grows both the planning horizon and the combinatorial complexity. While the planner (MCTD or a baseline) issues high-level positions for the cubes, a value-learning policy (Wang et al., 2023) executes local actions. We augment MCTD with object-wise guidance by selecting which cube is manipulated at each step, avoiding simultaneous movements of multiple cubes. We note that the object-wise guidance is particularly suited to MCTD because the tree-expansion mechanism naturally supports discrete ”meta-actions” at each node, such as selecting which object to move next. Single-pass diffusion methods, by contrast, attempt to produce a single, holistic trajectory in a single pass, leaving no opportunity to decide object-by-object. This guidance and additional task-specific reward shaping are described in Appendix A.12.
(a) Stacking Task
(b) Holistic Plan
(c) Object-wise Plan
Figure 4: (a) Illustration of the robot arm cube manipulation task. (b) A holistic plan for moving two cubes, where entanglement between their partial trajectories leads to an inaccurate movement (blue cube). (c) An object-wise plan produced by MCTD-Replanning, which avoids entanglement and achieves more reliable control over each cube’s movement.
Results
Table 2 shows success rates for one to four cubes. All methods perform comparably on single-cube tasks, but performance drops significantly for multi-cube scenarios, especially for standard diffusion baselines that struggle to decide cube ordering and handle subtasks such as stacking. However, MCTD-Replanning shows significant performance gaps from other models for multi-cube scenarios. MCTD exhibits moderate gains (e.g., 22% on two cubes), though it suffers from holistic plan entanglements when multiple objects are involved as shown in Figure 4b. To address this, MCTD-Replanning periodically replans, effectively separating each cube’s movements as shown in Figure 4c. This improves the success rate on the two-cube problem from 22% to 50%, and also benefits tasks with more cubes. Another challenge arises from the value-learning policy, which generalizes poorly as the number of cubes increases. Even if MCTD generates a suitable global plan, flawed local control undermines the final outcome. Interestingly, training the policy solely on single-cube data still allows MCTD-Replanning to solve a portion (24%) of four-cube scenarios, illustrating the potential of object-wise guidance and iterative replanning.
@@ 5.4 Visual Pointmaze
To evaluate our method on image-based planning, particularly under partial observability, we introduce a visual pointmaze environment. In this environment, the agent receives top-down pixel images of the initial and goal states, as shown in Figure 5. Unlike the visual antmaze in OGBench, which involves complex locomotion, this environment allows us to focus on how well the planner operates with raw visual inputs without relying on ground-truth states.
@@
We pre-encode observations with a Variational Autoencoder (VAE) (Kingma & Welling, 2014) to generate latent states suitable for diffusion-based planning, $z_{t}=\text{VAE}(x_{t})$. Because Diffuser’s default heuristic controller is inapplicable to purely visual inputs, we design an inverse dynamics model that maps consecutive latent observations to actions, $\hat{a}_{t}=f_{\text{inv}}(z_{t-1},z_{t},z_{t+1})$. For guidance, we employ a position estimator trained on ground-truth coordinates; however, this estimator does not update the underlying VAE, ensuring that the planner remains in a partially observable regime. Further details on this task are provided in Appendix A.13.
Results
@@
Table 3 compares success rates on medium and large versions of the visual pointmaze. MCTD and MCTD-Replanning show better performance than baselines on image-based, partially observable tasks. MCTD outperforms Diffuser baselines and Diffusion Forcing in the medium maze, presumably due to its tree-structured exploration, which better balances exploration and exploitation under perceptual partial observability. In the large maze, single-pass planners fail almost entirely, highlighting the difficulty of long-horizon visual planning. MCTD-Replanning improves performance by reevaluating partial plans mid-trajectory, though success remains modest overall. This result underscores the inherent challenge of scaling diffusion-based methods to extended horizons with visual data under partial observability and suggests that future work may need more sophisticated visual encoders or hierarchical guidance strategies.
@@
Figure 6: Success rates (%) and wall-clock runtime (Sec.) as the maximum number of denoising steps (test-time budget) increases for large and giant pointmaze tasks
5.5 Inference-Time Scalability & Time Complexity
@@
We examine how each planner leverages additional test-time computation by varying the maximum number of denoising steps and measuring both success rate and runtime on the large and giant pointmaze tasks (Figure 6). As the inference-time budget grows, MCTD consistently achieves high success rates—ultimately nearing perfection in giant mazes. In contrast, simply increasing denoising steps (Diffuser) or drawing more random samples (Diffuser-Random Search) yields modest returns, suggesting limited test-time scalability.
@@
MCTD incurs additional overhead from maintaining and expanding a tree—leading to a larger runtime as the budget grows. For instance, Diffuser-Random Search shows moderate runtime growth because multiple samples can be denoised in parallel batches, which is not directly feasible in MCTD’s tree-structured approach. Nonetheless, MCTD can terminate early whenever a successful plan is found, thereby limiting unnecessary expansions. As a result, on the large maze, MCTD’s runtime quickly saturates once the task can be solved with a smaller search depth, rather than fully consuming the available budget. A similar pattern emerges in the giant maze: as success rates increase under higher budgets, the runtime growth for MCTD remains sublinear, indicating that many expansions terminate early once suitable trajectories are located.
5.6 Ablation Studies
In this section, we conduct empirical ablation studies to analyze key components of MCTD, including the statistical nature of tree search, exploration-exploitation balance, tree depth, and the effectiveness of causal denoising and tree search. These ablation studies are primarily conducted on the PointMaze environment using the navigate datasets.
5.6.1 Greedy Tree Search
To evaluate the effectiveness of statistical tree search in balancing exploration and exploitation, we implemented a greedy tree search baseline. This approach searches candidates at each branching point and selects the best sample using the same reward function employed in Diffuser-Random Search and MCTD. For comprehensive evaluation, we varied the number of children for this approach from 5 to 20 (MCTD uses 5 children by default).
Table 4: Greedy Tree Search Performances (%) on PointMaze.
Children# 5 10 15 20
Medium 62$\pm$6 58$\pm$11 60$\pm$0 60$\pm$0
Large 40$\pm$0 40$\pm$0 40$\pm$0 40$\pm$0
Giant 0$\pm$0 0$\pm$0 0$\pm$0 0$\pm$0
The limited performance shown in Table 4 can be attributed to the method’s lack of deep search across multiple trajectory samples and the absence of backtracking mechanisms inherent in proper tree structures.
5.6.2 Meta-Action Set Investigation
A core mechanism of MCTD is balancing exploration and exploitation to generate globally optimal plans. In this section, we study this balance by analyzing different meta-action set configurations. We evaluate five meta-action sets ranging from exploration-biased to exploitation-biased configurations. Set 1 is exploration-biased, consisting of small guidance levels $[0,0.02,0.05,0.07,0.1]$, while Set 2 represents our default configuration $[0,0.1,0.5,1,2]$. Sets 3 and 4 are increasingly exploitation-biased with $[0.5,1,2,3,4]$ and $[4,5,6,7,8]$, respectively. Finally, Set 5 includes extreme guidance levels $[0.1,0.1,1,10,100]$ to test whether MCTD can effectively select appropriate guidance levels while ignoring extreme values.
Table 5: Performances of Diverse Meta-Action Sets on PointMaze.
Sets Set 1 Set 2 Set 3 Set 4 Set 5
Medium 88$\pm$16 100$\pm$0 100$\pm$0 98$\pm$6 100$\pm$0
Large 44$\pm$15 98$\pm$6 90$\pm$0 80$\pm$0 100$\pm$0
Giant 54$\pm$23 100$\pm$0 100$\pm$0 74$\pm$18 100$\pm$0
The results in Table 5 demonstrate that performance degrades when meta-actions are heavily biased toward either exploration (Set 1) or exploitation (Set 4). Notably, MCTD maintains robust performance with balanced meta-action sets, even when some values are extreme outliers (Sets 2, 3, and 5). This indicates that maintaining diverse guidance options is more critical than their specific values, provided both exploration and exploitation are adequately represented.
5.6.3 Subplan Length
Another key configurable component in MCTD is tree depth, which can be controlled through the subplan count $S$. While $S$ is adaptively chosen based on the environment’s episode length, in this section we investigate the impact of varying tree depths by manually controlling $S$.
Table 6: Empirical Results for Subplan Counts on PointMaze-Large.
$S$ 1 3 5 20
Success Rate 80$\pm$0 90$\pm$10 100$\pm$0 92$\pm$10
Run Time (sec.) 11$\pm$0 19$\pm$4 65$\pm$11 131$\pm$10
Search # 111$\pm$31 64$\pm$50 190$\pm$37 500$\pm$0
The results in Table 6 reveal an important trade-off between computational efficiency and performance. With longer subplans (smaller $S$), the search space is reduced, leading to faster execution times but potentially compromising performance quality. Conversely, very short subplans (e.g., $S=20$) require more search iterations, which may exhaust the computational budget before finding optimal solutions.
5.6.4 Causal Denoising and Tree Search
To investigate the effectiveness of causal denoising and tree search, we compared four variants: Diffusion Forcing without Causal Denoising (DFwoCD), Diffusion Forcing (DF), MCTD without Causal Denoising (MCTDwoCD), and MCTD.
Table 7: Performances of Causal Denoising and Tree Search Ablation on PointMaze.
Method DFwoCD DF MCTDwoCD MCTD
Medium 44$\pm$22 40$\pm$21 87$\pm$16 100$\pm$0
Large 50$\pm$14 55$\pm$20 78$\pm$6 98$\pm$6
Giant 8$\pm$10 34$\pm$16 18$\pm$11 100$\pm$0
The results reveal that both components contribute substantially to MCTD’s performance. Tree search consistently improves performance across all maze sizes. Meanwhile, causal denoising provides particularly substantial gains for long-horizon planning, with performance improvements most pronounced in the giant maze.
@@ 6 Limitations & Discussion
While MCTD enhances inference-time scalability by integrating search-based planning with diffusion models, it remains computationally expensive due to its System 2-style deliberative reasoning. Determining when to engage in structured planning versus relying on fast System 1 (model-free) planning is an open challenge, as expensive tree search may not always be necessary. Adaptive compute allocation based on task complexity or uncertainty could improve efficiency.
Another limitation is inefficiency in large-scale search spaces, where evaluating multiple trajectory hypotheses remains computationally demanding despite using low-dimensional meta-actions. A promising direction is amortized search, where the system meta-learns from inference-time search to improve exploration efficiency over time. Instead of treating initial exploration as random, MCTD could incorporate inference-time learning mechanisms to refine its search dynamically. Additionally, self-supervised reward shaping could improve trajectory evaluation in sparse-reward settings. Optimizing parallelized denoising, differentiable tree search, and model-based rollouts could further enhance efficiency.
@@ 7 Conclusion
We introduced Monte Carlo Tree Diffusion (MCTD), a framework designed to combine the best of both worlds: the structured search of Monte Carlo Tree Search and the generative flexibility of diffusion planning to enhance the inference-time scalability of System 2 planning. MCTD leverages meta-actions for adaptive exploration-exploitation, tree-structured denoising for efficient diffusion-based expansion, and fast jumpy denoising for rapid simulation. Experimental results demonstrate that MCTD outperforms existing approaches in various planning tasks, achieving superior scalability and solution quality. Future work will explore adaptive compute allocation, learned meta-action selection, and reward shaping to further enhance performance, paving the way for more scalable and flexible System 2 planning.
Acknowledgements
This research was supported by GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165) and Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT.
@@ A.1 Baselines
We compare MCTD against several baselines, each highlighting different ways to leverage inference-time scaling on diffusion models or broader planning strategies:
  • Diffuser (Janner et al., 2022): We use the standard single-shot Diffuser as our primary diffusion baseline. It generates a full trajectory by denoising from maximum to zero noise in one pass, guided by a return-predicting function $J_{\phi}$. While it can produce coherent plans, it lacks any mechanism to adapt or refine these plans at test time.
  • Diffuser-Replanning: To partially address the lack of iterative search, we evaluated a replanning variant of Diffuser. After an initial denoised plan is sampled, the policy is periodically re-invoked at predefined intervals (e.g., every 10 time steps), using the agent’s new state as the “start” for a new diffusion process. This strategy leverages additional test-time compute by attempting incremental corrections but does not share information across different replans.
  • Diffuser-Random Search: We consider a random-search variant of Diffuser that generates multiple trajectories from random noise and selects the best candidate according to a heuristic or learned value function. This approach can be seen as a “Sample-Score-Rank” method proposed in (Zhou et al., 2024), increasing test-time compute by drawing more samples but lacking the systematic exploration or adaptation found in tree-search paradigms.
  • Diffusion Forcing: Diffusion Forcing (Chen et al., 2024a) extends Diffuser by introducing a tokenized, causal noise schedule. Sub-plans of the trajectory can be denoised at different rates, allowing partial re-sampling of specific time segments without re-optimizing the entire plan. In our experiments, we evaluate Diffusion Forcing with the Transformer backbone provided in their official repository, which performed better in our preliminary tests. As the Diffusion Forcing baseline is inherently designed for replanning, we follow that evaluation setting.
A.2 Heuristic Controller, Value-Learning Policy, and Inverse Dynamics Model
A long-standing challenge in diffusion-based planning is preserving global trajectory coherence while managing local, high-dimensional state-action control (Chen et al., 2024a, b). For example, PlanDQ (Chen et al., 2024b) couples a high-level diffusion planner with a learned low-level policy. Similarly, our approach focuses the diffusion planner on lower-dimensional, representative state information (e.g., object or agent positions), while delegating detailed action inference to a low-level action module (i.e., a heuristic controller, a value-learning policy, or an inverse dynamics model). The details of this integration are discussed in Algorithm 10. Such a hierarchical design allows MCTD and baseline planners to concentrate on broader strategic decisions while leaving the finer-grained execution details to this action module.